a 13.4g sample of an unknown liquid is vaporized at 85degrees C and 100.0kPa. The vapour has a volume of 2.85L. The percentage composition of the liquid is found to be 68.5% carbon 8.6% hydrogen and 22.8% oxygen. What is the molecular formula of the liquid? Please show all steps! Thanks !!!
1
Expert's answer
2012-06-19T09:16:27-0400
pV=nRT
n = pV/RT = 100.0 kPa * 2.85 L / (358 K)*(8.314 L*kPa*K−1*mol−1) =
0.096 mol
M = m/n = 13.4 g / 0.096 mol = 139.6 g/mol
Determination of the
simplest moles ratio:
CxHyOz
x:y:z = 68.5/12 : 8.6/1 : 22.8/16 = 5.71 :
8.6 : 1.425 = 4 : 6 : 1
empirical formula:
M(C4H6O) = 4*12 + 6 + 16 = 70
g/mol
Calculating the common factor:
The common factor defines the ratio
(molecular formula)/(empirical formula):
139.6/70 = 1.99 = 2
Determination
of the molecular formula :
The solution: Multiply C4H6O by the common
factor -->
the true formula is (C4H6O)2 or C8H12O2